OpenStudy (anonymous):

can someone explain to me how to add 5/20 and 6/15 I don't understand lowest common denominators?

OpenStudy (anonymous):

A really easy way to do this (although not the most efficient) is to just start listing all the multiples of the denominators:

OpenStudy (julie):

Yeah, basically you can't add two fractions unless they have the same denominator.

OpenStudy (anonymous):

the denoms are 20 and 15, so list the multiples: 15: 15, 30, 45, 60, 75 ... 20: 20, 40, 60, 80 ...

OpenStudy (anonymous):

here you can move from left to right until you find one that's in both lists :)

OpenStudy (anonymous):

hmmm, ok, so how would I add them?

OpenStudy (julie):

misterf nailed it :)

OpenStudy (julie):

Ok so once you find the common denominator -- above, we see that it would be 60.

OpenStudy (siddharth):

Well, to add or subtract fractions, you have to make sure the bottom numbers on both fractions are the same. So, the best way to add two fractions with different denominators is to find the Lowest Common Multiple, and change the fractions to that form to perform the addition.

OpenStudy (julie):

Then you find how you make the denominators of both the same as that common one. So 15 * 4 = 60 and 20 * 3 = 60.

OpenStudy (julie):

Then you multiply the numerator, too: 6 * 4 = 24 15 * 4 = 60 and 5 * 3 = 15 20 * 3 = 60

OpenStudy (anonymous):

\[\frac{5}{20} = \frac{15}{60}\] \[\frac{6}{15} = \frac{24}{60}\] then add them normally: \[\frac{15}{60} + \frac{24}{60} = \frac{39}{60}\] does this make sense?

OpenStudy (julie):

Haha misterf beat me to it :)

OpenStudy (anonymous):

please gimme a medal if i helped i am collecting them :p

OpenStudy (julie):

When you add two fractions and they have the same denominator, the denominator stays the same and you just add the numerators.

OpenStudy (julie):

Gave it! :)

OpenStudy (siddharth):

Misterf explained it really nicely. Let us know if you don't understand something, albertw. I gave you a medal as well, misterf :)

OpenStudy (anonymous):

yay!!